A thin flake of transparent material (n = 1.44) is used to cover one slit of a double-slit interference arrangement. The central point on the screen is now occupied by what had been the 3th bright side fringe (m = 3). If λ = 574 nm, what is the thickness of the flake in meters?
Answer #1

in a double slit normally the path difference = d sin

in this case 1 slit is covered by snow of thickness t (say), due to this additional path difference = t (n-1)

total path difference = d sin + t x (n-1)

for location of original central maximum , =0

so at = 0 , total path difference = 0+ t (n -1)= m x = 3 x 574 x 10^(-9) m ( since m=3)

t = 3 x 574 x 10^(-9) m / (1.44 - 1)= 3.91 micro meters

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image

We were unable to transcribe this image